father is 32 years older than the daughter. in 4 years he will be 5 times her age. What is the age of the father and daughter

Answer by JulietG(1812)

Let x be the daughter's age.
We know that Dad is x + 32.
.
Jump ahead 4 years.
Now the daughter's age is x+4, and the Dad's age is x+36.
We're told that at this time, he is 5 times her age, so
5(x+4) = x+36
Distribute
5x+20 = x+36
Subtract 20 from each side:
5x = x+16
Subtract x from each side:
4x = 16
Divide by 4 on each side:
x = 4
If the daughter is 4, then Dad must be 36 (4+32).
.
Let's check it to make certain by looking 4 years ahead. If the calculations are correct:
In 4 years, the daughter will be 8.
In 4 years, Dad will be 40.
40 is indeed 5 times 8.